OpenStudy (perl):

Find the zeroes algebraically f(x) = 2x^4 -2x^2 -40 I dont understand the book answer

OpenStudy (perl):

please explain step by step

OpenStudy (perl):

the book got + - sqrt 5

OpenStudy (campbell_st):

well start by taking 2 as a common factor \[f(x) = 2(x^4 - x^2 -20)\] you have an equation that can be factorised \[f(x) = 2(x^2 + 4)(x^2 - 5)\] so you need to solve \[x^2 + 4 = 0\] and \[x^2 - 5 = 0\]

OpenStudy (perl):

i dont get how you factorised it

OpenStudy (perl):

ok let me try this, u = x^2

OpenStudy (perl):

so 2 ( u^2 - u-20)

OpenStudy (campbell_st):

thats it... so it factors to 2(u + 4)(u - 5) = 0

OpenStudy (anonymous):

Let's first factor out the 2: \(f(x)=2(x^4-x^2-20) = 0 \implies x^4-x^2-20=0\). Now let \(u=x^2\). So now we get: \(u^2-u-20=0\), now solve this quadratic normally for u and after getting the solutions, plug \(x^2\) back in for \(u\). @perl

OpenStudy (campbell_st):

I love it when people rewrite a previous post to explain things...

OpenStudy (anonymous):

@campbell_st I had written 90% of that when you posted yours.

OpenStudy (perl):

thankyou genius

OpenStudy (perl):

you guys are both great

OpenStudy (anonymous):

So stop loving what I wrote because this is my own solution, not copied off yours to show that I know how to do things.

OpenStudy (perl):

but wait, my book says the zeroes are sqrt(5) and -sqrt(5) ,

OpenStudy (perl):

they ignored the 2i and -2i

OpenStudy (campbell_st):

thats correct as the other possible zeros are complex x^2 = -4... is undefined in the real number system

OpenStudy (perl):

so by 'zeroes' they mean real zeroes?

OpenStudy (campbell_st):

yes... as you are graphing on a number plane of real numbers...

OpenStudy (perl):

hmmm

OpenStudy (anonymous):

After solving for u, you get: \((u-5)(u+4)=0 \implies u=5,-4 \) Plug \(x^2\) back in for u and solve for x: \(x^2=5 \implies (x-5)(x+5)=0 \implies x = \pm 5.\) \(x^2=-4 \implies x^2+4=0, x \in \emptyset\) \(\therefore x = \pm 5\) Do you understand now? @perl

OpenStudy (perl):

then when do we include *all* zeroes, all roots?

OpenStudy (anonymous):

Normally, questions only ask for real zeroes. You don't give the complex zeroes unless asked to. @perl

OpenStudy (perl):

ok thanks :)))

OpenStudy (perl):

y = 4x^3 -20x^2+25x

OpenStudy (perl):

find the real zeroes

OpenStudy (campbell_st):

well take x out as a common factor and you have \[y = x(4x^2 + 20x + 25)\] the quadratic in brackets is a perfect square... so you just need to factor the quadratic for a repeated zero... by taking out x as a common factor you have a zero at x = 0 hope this helps